# William W. Cohen1, Haohan Wang4, Jiaxin Huang2 1CMU, 2WUSTL, 3UW, 4UIUC

### v1

@@ Abstract
Existing reasoning evaluation frameworks for Large Language Models (LLMs) and Large Vision-Language Models (LVLMs) predominantly either assess text-based reasoning or vision-language understanding capabilities, with limited dynamic interplay between textual and visual constraints. To address this limitation, we introduce CrossWordBench, a benchmark designed to evaluate the reasoning capabilities of both LLMs and LVLMs through the medium of crossword puzzles—a task requiring multimodal adherence to semantic constraints from text-based clues and intersectional constraints from visual grid structures. CrossWordBench leverages a controllable puzzle generation framework that produces puzzles in multiple formats (text and image) and offers different evaluation strategies ranging from direct puzzle solving to interactive modes. Our extensive evaluation of over 20 models reveals that reasoning LLMs outperform non-reasoning models substantially by effectively leveraging crossing-letter constraints. We further demonstrate that LVLMs struggle with the task, showing a strong correlation between their puzzle-solving performance and grid-parsing accuracy. Our findings offer insights into the limitations of the reasoning capabilities of current LLMs and LVLMs, and provide an effective approach for creating multimodal constrained tasks for future evaluations.
@@ 1 Introduction
Figure 1: Performance gap between top-performing reasoning LLMs and non-reasoning LLMs/LVLMs on CrossWordBench. Reasoning LLMs achieve better overall performance vs. non-reasoning models, and follow crossing letter constraints significantly better.
@@
Large reasoning models (e.g., OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025)) have made exceptional progress on reasoning benchmarks including math problem solving (Veeraboina, 2023), coding (Austin et al., 2021) and commonsense reasoning (Clark et al., 2018). However, existing evaluation benchmarks for Large Language Models (LLMs) and Large Vision-Language Models (LVLMs) (e.g., Visual Question Answering (Yue et al., 2023)) mostly focus on text-based reasoning or vision-language understanding, lacking the dynamic interplay between textual and visual constraints that characterizes real-world problem solving. Consequently, evaluating multimodal reasoning capability—particularly for reasoning tasks requiring both textual and visual constraints—remains challenging.
@@
Crossword puzzles, a classic grid-based task in which horizontal (“Across”) words and vertical (“Down”) words must be filled in based on text-based clues, provides a unique testbed for such evaluations. They pose two distinct challenges: (1) question answering for each text-based clue, which may admit multiple correct solutions, and (2) visual constraint satisfaction, which requires precise letter alignment at the intersections of Across and Down entries. Prior crossword datasets (Efrat et al., 2021; Rozner et al., 2021; Kulshreshtha et al., 2022; Chen et al., 2025) often suffer from their reliance on static and copyrighted online news sources and adopt a text-centric formulation, thereby neglecting the visual structure.
@@
In this paper, we introduce CrossWordBench, a controllable and scalable benchmark for evaluating the reasoning capabilities of both LLMs and LVLMs. CrossWordBench collects data and generates puzzles from three sources: (1) multilingual word-clue pairs from public repositories, (2) dictionary-based definitions, and (3) adapted questions-answer pairs from existing benchmarks (e.g., CommonsenseQA (Talmor et al., 2018)) where the answers are open-ended or unconstrained. By representing grids and clues in different formats, as shown in Figure 2, CrossWordBench facilitates the evaluation of both model types. It supports two evaluation modes: a direct puzzle-solving mode, which generates one-round responses using zero-shot Chain-of-Thought (CoT) prompts, and an interactive mode for step-by-step puzzle-solving, where a grid generator is called to provide intermediate visual outputs for follow-ups, thereby serving as a foundation for evaluating agents using function calling.
@@
We evaluate over 20 state-of-the-art models, including both proprietary models (e.g., GPT-4o (Hurst et al., 2024) and Claude 3.7 Sonnet (Anthropic, 2024)) and open-weight models (e.g., DeepSeek-R1 (Guo et al., 2025) and Pixtral-Large-Instruct (Agrawal et al., 2024)). Our evaluation yields several notable findings: (1) LVLMs perform significantly worse than LLMs on CrossWordBench (as shown in Figure. 1), and they struggle in OCR in vertical (“Down”) word extraction. In fact, their puzzle-solving performance strongly correlates with their grid-parsing accuracy ($r=0.94$). (2) Reasoning LLMs outperform non-reasoning models, and benefit from test-time scaling and increased crossing-letter constraints. (3) Even puzzles derived from saturated benchmarks (e.g., CommonsenseQA) remain challenging, highlighting the significance of structural constraints in reasoning evaluation.
@@ 2 Related Work
LLMs and LVLMs Reasoning.
Recent research on the reasoning capabilities of LLMs (Ahn et al., 2024; Huang & Chang, 2022; Kojima et al., 2022; Plaat et al., 2024; Jaech et al., 2024; Guo et al., 2025; OpenAI, 2025; Huang et al., 2022) has led to the development of various approaches, including prompting-based methods (Wei et al., 2022; Yao et al., 2023; Besta et al., 2024; Chen et al., 2022) that guide LLMs to generate intermediate steps for solving complex problems, fine-tuning methods that train LLMs on long reasoning chains (Ye et al., 2025; Muennighoff et al., 2025; Zhao et al., 2025), and test-time scaling via self-refinement or with a verifier (Setlur et al., 2024; Feng et al., 2023; Wang et al., 2023; Zhang et al., 2024a; Uesato et al., 2022; Huang et al., 2025a). Moreover, a recent study from Deepseek-R1 (Guo et al., 2025) demonstrates that reinforcement learning (RL) with verifiable rewards facilitates the emergence of complex thinking processes in LLMs. Several studies have explored reasoning in the multimodal domain by constructing CoT data for fine-tuning (Xu et al., 2024) and developing verifiable problems for RL (Yang et al., 2025; Huang et al., 2025b). These approaches have demonstrated success, yet the evaluation datasets remain largely restricted to mathematical problems (Wang et al., 2024; Lu et al., 2023; Zhang et al., 2024b).
Crossword Puzzles in Language Model Evaluation.
Crossword puzzles have long been a focus of research in natural language processing (NLP), particularly before the advent of LLMs. Early approaches typically employed constraint satisfaction algorithms augmented by external knowledge bases. Notable examples include systems such as Proverb (Littman et al., 2002), Dr. Fill (Ginsberg, 2011), and specialized models such as the Berkeley Crossword Solver (Wallace et al., 2022), which incorporate a fine-tuned BERT (Devlin et al., 2019) and belief propagation. More recent studies have leveraged LLMs to address crossword puzzles through techniques including fine-tuning (Efrat et al., 2021; Rozner et al., 2021; Sadallah et al., 2024), prompting strategies such as Tree-of-Thoughts (Yao et al., 2023), or integration with search algorithms (Saha et al., 2024), demonstrating the potential of LLMs for crosswords.
@@ 3 Benchmark Curation
Online sources, including The New York Times and The Los Angeles Times, provide complete crossword puzzles. Nevertheless, directly utilizing these puzzles presents several disadvantages. (1) Copyright restrictions may limit their usage, and their online availability increases the probability that they have been incorporated into the pretraining datasets of contemporary models. (2) Online puzzles are typically static—with predefined words, themes, grid sizes, and strict formatting (e.g., black square symmetry) [^1] The New York Times Crossword Requirements.—which not only limits their adaptability for diverse benchmark generation, but also imposes arbitrary formatting constraints that do not yield meaningful benefits for evaluation. Prior research (Mirzadeh et al., 2024) has shown that even minor modifications to questions can significantly affect model performance, particularly when performances on the original version is near saturation. (3) The static characteristics of these puzzles restricts the range of potential evaluation strategies. To overcome these limitations, we propose a two-stage benchmark construction strategy:
@@
While the possibility of data contamination may still exist for these word-clue pairs, the variations in grid design can yield distinct crossword puzzles, and as demonstrated in Section 4.5, even when incorporating extensively tuned data such as CommonsenseQA training set, the resulting puzzles remain remarkably challenging for both LLMs and LVLMs.
@@
Total ##\## of puzzles 100 100 100 100 50
Total ##\## of words 1,193 3,472 1,327 1,139 543
Unique words (total) 83.82% 80.76% 92.92% 36.70% 59.85%
Unique clues (total) 100% 100% 100% 100% 100%
@@
Table 1: Crossword puzzle statistics for different subjects and grid sizes. Statistics are presented separately for each category, as distinct word and clue pairs are used for their construction. Additionally, aggregated statistics for all English puzzles set are included.
Puzzle Generation. After collecting word-clue pairs, we implement an automatic pipeline to generate puzzles in both text and image formats. The pipeline enables (1) puzzle difficulty control by adjusting grid sizes and (2) incremental word placement function, which facilitates an interactive evaluation mode described in Section 4.4. Figure 2 illustrates the overall benchmark pipeline. To maintain category integrity, we avoid overlapping clues under each category. Table 1 presents the overall dataset statistics for each category.
@@ 4 Experiment
In the following sections, we present an extensive empirical evaluation of CrossWordBench across multiple model architectures and examine their performance characteristics.
@@ 4.1 Experimental Setup
Evaluation Data and Metrics.
For main experiments, we evaluate models on the English set with three metrics to assess answer accuracy and adherence to crossword constraints.
@@
Evaluated Models.
We evaluate a collection of proprietary and open-weight models, including both LVLMs and LLMs. For proprietary models, we consider state-of-the-art models such as GPT-4o (Hurst et al., 2024) and Claude-3-7-Sonnet (Anthropic, 2024) (with and without thinking mode). For open-weight models, our selections ranging from 3B to 124B parameters, such as the Qwen Series (Team, 2024; Bai et al., 2025) and Pixtral-Large-Instruct-2411 (Agrawal et al., 2024) for LVLMs. For LLMs, we evaluate both reasoning models such as the Deepseek Series (Liu et al., 2024; Guo et al., 2025) and non-reasoning models such as Llama series (Dubey et al., 2024). The full list of models is shown in Table 2. We set temperature to 0 for non-reasoning models for consistency, 0.6 for reasoning models based on commonly recommended settings, and 1.0 to certain proprietary models (e.g., used by Claude-3-7-Sonnet Thinking). Further generation details are in Appendix B.3.
Input Prompt Templates and Output Response Parsing.
For the main evaluation, we adopt the zero-shot Chain-of-Thought (CoT) (Wei et al., 2022) prompting strategy. In LVLM evaluation, both the clues and the grid are embedded within an image. In LLM evaluation, the grid is represented as a 2D binary array, with 1 indicating a blocked cell and 0 representing an unfilled cell, and is preprended to text clues in the prompt. Detailed prompt templates are in Appendix B.4. To extract answers from responses, we leverage the structured output capabilities of o3-mini [^5] https://openai.com/index/openai-o3-mini/ to convert raw model responses into JSON format, by generating dynamic Pydantic models. Implementation details are in Appendix B.2.
@@ 4.2 Main Results
[…]
Reasoning LLMs substantially outperform conventional ones across metrics, with notable improvements in ICR, as shown in Table 2. In particular, among reasoning LLMs, o3-mini achieves an ICR of 0.891 on 7x7 grid, demonstrating a strong ability to interpret and enforce grid constraints. Although DeepSeek-R1 attains the highest WCR and LCR on 7x7 grid, its ICR is slightly lower than that of o3-mini. Other reasoning models, such as QwQ-32B and R1-Distilled-Llama-70B, show moderate performance, with WCRs of approximately 0.347 and 0.387, respectively. One notable outlier is Open-Reasoner-Zero-32B, which performs poorly across all metrics, indicating that it does not effectively leverage grid constraints for reasoning. This suggests that its training—primarily focused on mathematical reasoning—does not generalize well to tasks requiring spatial and linguistic integration, thereby highlighting a key limitation in the training strategies of these reasoning models.
Non-reasoning LLMs exhibit limitations, especially in ICR, and larger grids present increased challenges. Among non-reasoning LLMs, Claude-3-7-Sonnet and Gemini 2.0 Pro Exp yield the best results, with WCRs of 0.482 and 0.460 on the 7x7 grid, respectively; however, their relatively lower ICR indicates model limitations on explicit reasoning constraints for crossword puzzles. Notably, thinking mode improves Claude-3-7-Sonnet on all three metrics, highlighting the importance of reasoning and reflection in solving constraints in crossword tasks. Additionally, larger grids lead to decreasing performance, demonstrating the increased complexity of maintaining constraint adherence over a larger search space.
LVLMs currently lag behind LLMs in performance, with minimal adherence to grid constraints. With image inputs, Claude-3-7-Sonnet achieves the highest performance among LVLMs, but underperforms its own text inputs version. Models such as GPT-4o and Gemini 2.0 Pro Exp exhibit similar trends, with WCRs below 0.35 on larger grids. All LVLMs demonstrate low ICRs, suggesting that they struggle to maintain reasoning consistency. Notably, performance declines when thinking mode is enabled on Claude-3-7-Sonnet with image input, which contradicts the improvements observed with text inputs; we explore this phenomenon in the next section. Among open-weight LVLMs, Pixtral-Large-Instruct achieves the best WCR of 0.297 on 7x7 grid, while still lags behind most proprietary LVLMs.

### v2

@@ Abstract
Existing reasoning evaluation frameworks for Large Language Models (LLMs) and Large Vision-Language Models (LVLMs) predominantly assess either text-based reasoning or vision-language understanding capabilities, with limited dynamic interplay between textual and visual constraints. To address this limitation, we introduce CrossWordBench, a benchmark designed to evaluate the reasoning capabilities of both LLMs and LVLMs through the medium of crossword puzzles—a task requiring multimodal adherence to semantic constraints from text-based clues and intersectional constraints from visual grid structures. CrossWordBench leverages a controllable puzzle generation framework that produces puzzles in two formats (text and image), supports adjustable difficulty through prefill ratio control, and offers different evaluation strategies, ranging from direct puzzle solving to interactive modes suitable for agentic evaluation. Our extensive evaluation of over 20 models reveals that reasoning LLMs substantially outperform non-reasoning models by effectively leveraging crossing-letter constraints. We further demonstrate that LVLMs struggle with the task, showing a strong correlation between their puzzle-solving performance and grid-parsing accuracy. Our findings highlight the limitations of the reasoning capabilities of current LLMs and LVLMs, and provide an effective approach for creating multimodal constrained tasks for future evaluations.
@@ 1 Introduction
Figure 1: Performance gap between top-performing reasoning LLMs and non-reasoning LLMs/LVLMs on CrossWordBench. Reasoning LLMs achieve better overall performance and adhere more effectively to crossing-letter constraints than non-reasoning LLMs/LVLMs.
@@
Large reasoning models (e.g., OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025)) have made exceptional progress on reasoning benchmarks including math problem solving (Veeraboina, 2023), coding (Austin et al., 2021) and commonsense reasoning (Clark et al., 2018). However, existing evaluation benchmarks for Large Language Models (LLMs) and Large Vision-Language Models (LVLMs) (e.g., Visual Question Answering (Yue et al., 2023)) mostly focus on text-based reasoning or vision-language understanding, lacking dynamic interplay between textual and visual constraints that characterizes real-world problem solving. Consequently, evaluating multimodal reasoning capabilities—particularly on reasoning tasks requiring both textual and visual constraints—remains challenging.
@@
Crossword puzzles, a classic grid-based task in which horizontal (“Across”) words and vertical (“Down”) words must be filled in based on text-based clues, provide a unique testbed for such evaluations. They pose two distinct challenges: (1) question answering for each text-based clue, which may admit multiple correct solutions, and (2) visual constraint satisfaction, which requires precise letter alignment at the intersections of Across and Down entries. Prior crossword datasets (Efrat et al., 2021; Rozner et al., 2021; Kulshreshtha et al., 2022; Chen et al., 2025) often suffer from their reliance on static and copyrighted online news sources and adopt a text-centric formulation, thereby neglecting the visual structure.
@@
In this paper, we introduce CrossWordBench, a controllable and scalable benchmark for evaluating the reasoning capabilities of both LLMs and LVLMs. CrossWordBench collects data and generates puzzles from three sources: (1) multilingual word-clue pairs from public repositories, (2) dictionary-based definitions, and (3) adapted question-answer pairs from existing benchmarks (e.g., CommonsenseQA (Talmor et al., 2018)) where the answers are open-ended or unconstrained. By representing grids and clues in different formats, as shown in Figure 2, CrossWordBench facilitates the evaluation of both model types. It supports two evaluation modes: a direct puzzle-solving mode, which generates one-round responses using zero-shot Chain-of-Thought (CoT) prompts, and an interactive mode for step-by-step puzzle-solving, where grid update functions can provide intermediate visual outputs for follow-ups, thereby serving as a foundation for evaluating agents using function calling.
@@
We evaluate over 20 state-of-the-art models, including both proprietary models (e.g., GPT-4o (Hurst et al., 2024) and Claude 3.7 Sonnet (Anthropic, 2024)) and open-weight models (e.g., DeepSeek-R1 (Guo et al., 2025) and Pixtral-Large-Instruct (Agrawal et al., 2024)). Our evaluation yields several notable findings: (1) LVLMs perform significantly worse than LLMs on CrossWordBench (as shown in Figure 1), and they struggle in OCR in vertical (“Down”) word extraction. In fact, their puzzle-solving performance strongly correlates with their grid-parsing accuracy ($r=0.94$). (2) Reasoning LLMs outperform non-reasoning models, and benefit from both test-time scaling and increased crossing-letter constraints. (3) Even puzzles derived from saturated benchmarks (e.g., CommonsenseQA) remain challenging, highlighting the significance of structural constraints in reasoning evaluation.
@@ 2 Related Work
LLMs and LVLMs Reasoning. Recent research on the reasoning capabilities of LLMs (Ahn et al., 2024; Huang & Chang, 2022; Kojima et al., 2022; Plaat et al., 2024; Jaech et al., 2024; Guo et al., 2025; OpenAI, 2025; Huang et al., 2022) has led to the development of various approaches, including prompting-based methods (Wei et al., 2022; Yao et al., 2023; Besta et al., 2024; Chen et al., 2022) that guide LLMs through intermediate reasoning steps for solving complex problems, fine-tuning methods that train LLMs on long reasoning chains (Ye et al., 2025; Muennighoff et al., 2025; Zhao et al., 2025), and test-time scaling via self-refinement or the use of a verifier (Setlur et al., 2024; Feng et al., 2023; Wang et al., 2023; Zhang et al., 2024a; Uesato et al., 2022; Huang et al., 2025a). Moreover, a recent study from Deepseek-R1 (Guo et al., 2025) demonstrates that reinforcement learning (RL) with verifiable rewards facilitates the emergence of complex thinking processes in LLMs. Several studies have explored reasoning in the multimodal domain by constructing CoT data for fine-tuning (Xu et al., 2024) and developing verifiable problems for RL (Yang et al., 2025; Huang et al., 2025b). While these approaches have demonstrated success, existing evaluation datasets remain largely restricted to math problems (Wang et al., 2024; Lu et al., 2023; Zhang et al., 2024b).
Crossword Puzzles in Language Model Evaluation. Crossword puzzles have long been a focus of research in natural language processing (NLP), particularly before the advent of LLMs. Early approaches typically employed constraint satisfaction algorithms augmented by external knowledge bases. Notable examples include systems such as Proverb (Littman et al., 2002), Dr. Fill (Ginsberg, 2011), and specialized models such as the Berkeley Crossword Solver (Wallace et al., 2022), which incorporate a fine-tuned BERT (Devlin et al., 2019) and belief propagation. More recent studies have leveraged LLMs to address crossword puzzles through techniques including fine-tuning (Efrat et al., 2021; Rozner et al., 2021; Sadallah et al., 2024), prompting strategies such as Tree-of-Thoughts (Yao et al., 2023), or integration with search algorithms (Saha et al., 2024), demonstrating the potential of LLMs for crosswords.
@@ 3 Benchmark Curation
Online sources, including The New York Times and The Los Angeles Times, provide complete crossword puzzles. Nevertheless, directly utilizing these puzzles presents several disadvantages. (1) Copyright restrictions may limit their usage, and their online availability increases the probability that they have been incorporated into the pretraining datasets of contemporary models. (2) Online puzzles are typically static—with predefined words, themes, grid sizes, and strict formatting (e.g., black square symmetry) [^1] The New York Times Crossword Requirements.—which not only limits their adaptability for diverse benchmark generation, but also imposes arbitrary formatting constraints that do not yield meaningful benefits for evaluation. Prior research (Mirzadeh et al., 2024) has shown that even minor modifications to questions can significantly affect model performance, particularly when performance on the original version is near saturation. (3) The static characteristics of these puzzles restrict the range of potential evaluation strategies. To overcome these limitations, we propose a two-stage benchmark construction strategy:
@@
While the possibility of data contamination may still exist for these word-clue pairs, the variations in grid design can still yield distinct crossword puzzles, and as demonstrated in Section 4.5, even when incorporating extensively tuned data such as CommonsenseQA training set, the resulting puzzles remain remarkably challenging for both LLMs and LVLMs.
Table 1: Crossword puzzle statistics for different subjects and grid sizes. Statistics are presented separately for each category, as distinct word and clue pairs are used for their construction. Additionally, the aggregated statistics for all English puzzles are included.
@@
Total #\# of puzzles 100 100 100 100 50
Total #\# of words 1,193 3,472 1,327 1,139 543
Unique words (% total) 83.82% 80.76% 92.92% 36.70% 59.85%
Unique clues (% total) 100% 100% 100% 100% 100%
@@
Puzzle Generation. After collecting word-clue pairs, we implement an automatic pipeline to generate puzzles in both text and image formats, as illustrated in Figure 2. The pipeline enables (1) puzzle difficulty control by adjusting grid sizes, and (2) incremental word placement functions, which facilitate the interactive evaluation mode described in Section 4.4.
The generation framework leverages a heuristic scoring function that encourages desirable properties such as a high number of horizontal and vertical word intersections. Post-processing filters are applied to discard puzzles with low clue counts or high blocked cell ratios, thereby enhancing overall quality and structural coherence. These heuristics and filters, in combination with grid size adjustments, help control puzzle complexity. To preserve the category integrity, we ensure that clues do not overlap within each category [^5] Please check out our code for more implementation details..
Puzzle Quality. While CrossWordBench does not include human-created puzzles, its design is inspired by them. Table 1 reports overall dataset statistics—including clue count, average word length, and blocked cell ratio—which serve as proxies for puzzle quality and difficulty.
Compared human-created puzzles—for example, those used LR2Bench (Chen et al., 2025), which were sourced from outlets such as The Los Angeles Times and Vulture—our generated puzzles exhibit comparable difficulty when evaluated using similar word-level metrics. Our focus on automatically generated puzzles is motivated by the need for scalable, controllable, and diverse evaluation across subjects and formats. This enables systematic variation in difficulty through grid size, prefill ratio, and clue selection; supports step-by-step reasoning via an interaction mode; and facilitates fine-grained metric analysis beyond overall accuracy.
@@ 4 Experiment
In the following sections, we present an extensive empirical evaluation of CrossWordBench across multiple different model architectures and analyze their performance characteristics.
@@ 4.1 Experimental Setup
Evaluation Data and Metrics. For the main experiments, we evaluate models on the English set with three metrics to assess answer accuracy and adherence to crossword constraints.
@@
Evaluated Models. We evaluate a collection of proprietary and open-weight models, including both LVLMs and LLMs. For proprietary models, we consider state-of-the-art models such as GPT-4o (Hurst et al., 2024) and Claude 3.7 Sonnet (Anthropic, 2024) (with and without thinking mode). For open-weight models, our selections range from 3B to 124B parameters, such as the Qwen Series (Team, 2024; Bai et al., 2025) and Pixtral-Large-Instruct-2411 (Agrawal et al., 2024) for LVLMs. For LLMs, we include both reasoning models such as the Deepseek Series (Liu et al., 2024; Guo et al., 2025) and non-reasoning models such as Llama series (Dubey et al., 2024). The full list of models is shown in Table 2. We set decoding temperature to 0 for non-reasoning models for consistency, 0.6 for reasoning models based on commonly recommended settings, and 1.0 for certain proprietary models (e.g., used by Claude 3.7-Sonnet with Thinking). Further generation details are provided in Appendix D.3.
Input Prompt Templates and Output Response Parsing. For the main evaluation, we adopt the zero-shot Chain-of-Thought (CoT) (Wei et al., 2022) prompting strategy. In LVLM evaluation, the clues and the grid are both embedded within an image. In LLM evaluation, the grid is represented as a 2D binary array, with 1 indicating a blocked cell and 0 representing an unfilled cell, and is prepended to text clues in the prompt. To extract answers from responses, we leverage the structured output capabilities of o3-mini [^6] https://openai.com/index/openai-o3-mini/ to convert raw model responses into JSON format by generating dynamic Pydantic models. Detailed prompt templates and implementation details are listed in Appendix D.4 and D.2.
@@ 4.2 Main Results
Reasoning LLMs substantially outperform conventional ones across metrics, with notable improvements in ICR, as shown in Table 2. In particular, among reasoning LLMs, o3-mini achieves an ICR of 0.891 on 7x7 grids, demonstrating a strong ability to interpret and enforce grid constraints. Although DeepSeek-R1 attains the highest WCR and LCR on 7x7 grids, its ICR is slightly lower than that of o3-mini. Other reasoning models, such as QwQ-32B and R1-Distilled-Llama-70B, show moderate performance, with WCRs of approximately 0.347 and 0.387, respectively. One notable outlier is Open-Reasoner-Zero-32B, which performs poorly across all metrics, indicating that it does not effectively leverage grid constraints for reasoning. This suggests that its training—primarily focused on mathematical reasoning—does not generalize well to tasks requiring spatial and linguistic integration, thereby highlighting a key limitation in the training strategies of these reasoning models.
Non-reasoning LLMs show limitations in ICR, and performance declines further with increasing grid size. Among non-reasoning LLMs, Claude 3.7 Sonnet and Gemini 2.0 Pro Exp yield the best results, with WCRs of 0.482 and 0.460 on the 7x7 grid, respectively; however, their relatively lower ICR indicates model limitations on explicit reasoning constraints for crossword puzzles. Notably, thinking mode improves Claude 3.7 Sonnet on all three metrics, highlighting the importance of reasoning and reflection in solving constraints-based crossword tasks. Additionally, larger grids lead to decreasing performance, demonstrating the increased complexity of maintaining constraint adherence over a larger search space.
LVLMs currently lag behind LLMs in performance, with minimal adherence to grid constraints. With image inputs, Claude 3.7 Sonnet achieves the highest performance among LVLMs, but underperforms its own text inputs version. Models such as GPT-4o and Gemini 2.0 Pro Exp exhibit similar trends, with WCRs below 0.35 on larger grids. All LVLMs demonstrate low ICRs, suggesting that they struggle to maintain reasoning consistency. Notably, performance declines when thinking mode is enabled on Claude 3.7 Sonnet with image input, which contradicts the improvements observed with text inputs; we explore this phenomenon in the next section. Among open-weight LVLMs, Pixtral-Large-Instruct achieves the best WCR of 0.297 on 7x7 grid, while still lags behind most proprietary LVLMs.